OpenStudy (anonymous):

multivariable calculus equation of line help: a) Find parametric equations for the line through (2,4,6) that is perpendicular to the plane x-y+3z=7 b) In what points does this line intersect the coordinate planes?

OpenStudy (turingtest):

the gradient of the equation of the plane will give you a vector \(\vec n\) perpendicular to the plane what do you have for the gradient?

OpenStudy (turingtest):

the parametric equation of a line requires a point, and a vector parallel to the line in this case we want a vector parallel to \(\vec n=\nabla f(x,y,z)\)

OpenStudy (anonymous):

we never learned gradiant yet...just basic vectors (dot product, cross product, etc)

OpenStudy (turingtest):

then choose three random points in the plane (1,0,2) (7,0,0) (0,-7,0) and get two vectors from those points what do you get for that?

OpenStudy (turingtest):

what vector can you get between (1,0,2) and (7,0,0) ?

OpenStudy (anonymous):

oh okay! so we just choose random points? (0,0,0) would work right?

OpenStudy (turingtest):

no, because (0,0,0) does not satisfy x-y+3z=7

OpenStudy (anonymous):

oh!!!

OpenStudy (turingtest):

\[0-3(0)+3(0)\neq7\]

OpenStudy (turingtest):

typo, but you get the point so choose some logical points by inspection

OpenStudy (anonymous):

so we get a point that satisfies that equation, then subtract it from the other point (2,4,6) to get a vector

OpenStudy (anonymous):

then use dot product?

OpenStudy (turingtest):

no, we are to going to use the given point just yet; that's for later

OpenStudy (turingtest):

aren't*

OpenStudy (turingtest):

first we need 3 points in the plane call them P1 P2 and P3 they must all be (x,y,z) that satisfy x-y+3z=7

OpenStudy (turingtest):

so, for instance, it looks like \((3,-1,1)\) is in the plane, because\[3-(-1)+3(1)=7\]

OpenStudy (anonymous):

okay

OpenStudy (turingtest):

so let\[P_1=(3,-1,1)\]\[P_2=(4,-2,0)\]\[P_3=(1,0,2)\]what vector do we get between\(P_1\) and \(P_2\) ?

OpenStudy (anonymous):

<1,-3,-1>?

OpenStudy (turingtest):

<1,-1,-1>

OpenStudy (anonymous):

btw my the points i chose were... (7,0,0) (0,-7,0) (4,0,1) but yeah i wna know the process.. my points are right?

OpenStudy (turingtest):

yeah, those work too

OpenStudy (anonymous):

oh yeah my bad lol -2-(-1) = -1 lol

OpenStudy (anonymous):

okay so after you get a vector...

OpenStudy (anonymous):

we get the vector P2P3?

OpenStudy (turingtest):

yes^ (thinking if sign matters; I think it does not)

OpenStudy (anonymous):

then you get the cross product of the two? since it's asking for a perpendicular plane?

OpenStudy (turingtest):

you got it :)

OpenStudy (anonymous):

oh and thats the answer for part A?! Thanks! and for part B we just set them equal to each other and solve for x,y,z right?

OpenStudy (turingtest):

whoa now, we're not done with A just yet we got the vector, but they want the \(line\) that passes through the given point, which we haven't used yet to do that you need to know that the parametric equation of a line parallel to a vector \(\vec v\) is given by\[\vec r(t)=P_0+t\vec v\]

OpenStudy (turingtest):

in our case \(\vec v=\vec n\) which is the cross product of the two vectors we just found, and \(P_0\) is the given point

OpenStudy (anonymous):

oh so we use P0 (which was given) +t(the vector we found from the 3 points)

OpenStudy (turingtest):

exactly

OpenStudy (anonymous):

I was wondering where the equation of a line concept came in, since it is equation of line section. Thanks!

OpenStudy (anonymous):

my I ask, how did you know right away to pick any points that satisfy x-y+3z=7? Like why would it make sense to use points that agree to that equation? Nothing in the problem gives it off for me.

OpenStudy (turingtest):

well I am familiar with the behavior of vectors in planes, and I know that if we pick 3 points in the same plane like so...

OpenStudy (anonymous):

oh it lies in the same plane...

OpenStudy (turingtest):

we can draw vectors between them